OpenStudy (anonymous):

The cost of printing a flyer to spread knowledge of an organization that you run is a flat rate of $9 plus $0.05 per copy. Let C represent the total cost of printing and x represent the number of copies. Write in symbolic form an equation that will relate C and x. Organize the data into a table of values. Begin with 500 copies, increase by increments of 500, and end with 2500 copies. How much would it cost to print 2500 flyers and what equation would you use?

OpenStudy (texaschic101):

9 + .05x = C 500 copies : 2000 copies 9 + .05x = C 9 + .05x = C 9 + .05(500) = C 9 + .05(2000) = C 9 + 25 = C 9 + 100 = C $ 34 = C 109 = C 1000 copies 2500 copies 9 + .05x = C 9 + .05x = C 9 + .05(1000) = C 9 + .05(2500) = C 9 + 50 = C 9 + 125 = C $ 59 = C 134 = C 1500 copies 9 + .05x = C 9 + .05(1500) = C 9 + 75 = C 84 = C your table..... x = 500 1000 1500 2000 2500 C = 34 59 84 109 134